what is The Ksp of AgCl at 25 oc is 1.0x10-10. Calculate the concentration of Ag+ and Cl- ions in a saturated solution of AgCl, and the molar solubility of AgCl ?
1
Expert's answer
2020-09-24T13:49:42-0400

AgCl <=> Ag+ + Cl-

[Ag+] = [Cl-]

Ksp = [Ag+][Cl-] = [Ag+]2 = 1.0×10-10

[Ag+] = 1.0×10-5 M

[Ag+] = [Cl-] = 1.0×10-5 M

Molar solubility = [Ag+] = 1.0×10-5 M